OpenStudy (anonymous):

Are the signs right? 4y^2-9z^2 (2y-9z)(2y-9z) F= 4y^2 O=-6yz I= -6zy L= 9z^2 =4y^2- 9z^2

OpenStudy (anonymous):

I am thinking maybe +- or ++ not - -

OpenStudy (anonymous):

(2y-9z)(2y-9z) F = 4y^2 O = -18yz I = -18yz L = 81z^2 Hence (2y-9z)(2y-9z) =4y^2-36yz+81z^2

OpenStudy (anonymous):

lol, I meant 3z not 9z.

OpenStudy (anonymous):

In that case... (2y-3z)(2y-3z) F = 4y^2 O = -6yz I = -6yz L = 9z^2 Hence (2y-3z)(2y-3z) =4y^2-12yz+9z^2

OpenStudy (anonymous):

it must be + + because +- wouldnt work. I dont think.